In Macroland, currency held by the public is 2,000 econs, bank reserves are 300 econs, and the desired (and current) reserve/deposit ratio is 15 percent. If commercial banks borrow 100 econs in reserves from the Central Bank through discount window lending, then the money supply in Macroland will ______, assuming that the public does not wish to change the amount of currency it holds.

Multiple Choice

  • increase to 4,100 econs

  • decrease to 1,900 econs

  • increase to 3,133 econs

  • increase to 4,667 econs
Answer #1

increase to 4,667 econs

(Original money supply = 2000 + (300/15%) = 2000 + (300/0.15) = 2000 + 2000 = 4000
New money supply = 2000 + [(300+100)/15%) = 2000 + (400/0.15) = 2000 + 2666.67 = 4666.67 = 4667 (approximately))